OpenStudy (anonymous):

Q3 The impulse response of a circuit is its response to a unit impulse, . Knowing the impulse response of a linear circuit is extremely valuable as we can figure out the circuit's response to an arbitrary input from it. In this problem you will find the response of a circuit when it is driven by a unit impulse.

OpenStudy (anonymous):

please give the answer i had not solved the 8th week home work

OpenStudy (anonymous):

Final MIT 6.002x Solution Q1 a)ia=2*i1*R/(2*R+rm) and vo=ia*rm b)iA =(-i2)*R/(2*R+rm) and V0=ia*rm c) part a) + part b) d)2R Q2 p1: Vs as vin<vt p2: v0 = vs-kr(e^((vin-vt)/vn)-1) p3: vin = ln(vs/2kr +1)vn+vt p4: vin=sqrt(vs/kr)+vt p5: -kre^((vin-vt)/vn) p6: -kr(vin-vt) Q3 same as HP8p1 Q4 p1=V*R2/(R1+R2)*(1-e^(-(R1+R2)*t/(R1*R2*C))) p2=(V*R2/(R1+R2))*cos(1/sqrt(L*C)*(t-T)) p3=C*(V*R2/(R1+R2))*sin((t-T)*1/sqrt(L*C))/sqrt(L*C) Q5 E V B Y F Z Q6 1) Time constant is L/Rtotal -> 1/5=0.2 2) Time constant is L/Rtotal -> 1/5=0.2 3) Vmax is 5.95840 V 4) Vmin is 0.040 V

OpenStudy (anonymous):

Beware of Q 6 values are different

OpenStudy (anonymous):

even if u havnt solved it u can view the answers..